OpenStudy (anonymous):

pleaaaase help D:

OpenStudy (anonymous):

Rewrite the rational exponent as a radical.

OpenStudy (anonymous):

help please... :(( the rational exponent is in the picture attached

OpenStudy (anonymous):

Just Multiplication powers ! means : 3/4 * 2/3 and get it for power .! we have : 5^3/4*2/3=5^2 And in the answers : We have :

OpenStudy (ranga):

\[(x ^{a})^{b} = x ^{ab}\] Identify x, a and b in your problem and substitute.

OpenStudy (anonymous):

okay well "x" is 5, "a" is 3/4 and "b" is 2/3 so it would be 5^ 3/4 2/3 ??

OpenStudy (ranga):

Yes!

OpenStudy (anonymous):

how is that translated into a radical?

OpenStudy (anonymous):

And in the radicals : \[\sqrt[y]{x^n}=x^n/y\]

OpenStudy (anonymous):

and we have the answers to try them : in the 1 : we have : 5^3/2 In 2 : 5^1/12 in 3 : 5^1/2 In 4 : 5^3/4 Whats we have in 5 ?

OpenStudy (anonymous):

@harbingers

OpenStudy (ranga):

\[5^{\frac{ 3 }{ 4 }\times \frac{ 2 }{ 3 }} = 5^{\frac{ 1 }{ 2}} = \sqrt{5}\]

OpenStudy (anonymous):

we should find a answer that if we divided that power on the radical power will be 1/2 .! I had a fall ! It s 1/2 not 2 .!

OpenStudy (anonymous):

so is the answer the third option down?? :O

OpenStudy (ranga):

Yep.

OpenStudy (anonymous):

omg okay thank you so much! imma give you a medal :)))

OpenStudy (ranga):

u r welcome!

OpenStudy (anonymous):

Yes !